A social club charters a plane at a cost of
$34848


to take a group of members on an excursion to Atlantic City.

At the last minute,
11
people in the group decide not to go.

This raises the transportation cost per person by
$24
.

How many people originally intended to take the trip?

Answer by stanbon(75887)

A social club charters a plane at a cost of
$34848 to take a group of members on an excursion to Atlantic City.
At the last minute, 11 people in the group decide not to go.
This raises the transportation cost per person by
$24.
How many people originally intended to take the trip?
----
Let original number be "x".
Average cost for the original number was 34848/x dollars.
---
New number of people is "x-11"
Average cost for the new number is 34848/(x-11) dollars.
--------------------
Equation:
new average - old average = 24 dollars
---
34848/(x-11) - 34848/x) = 24
----
34848x - 34848(x-11) = 24x(x-11)
+11*34848 = 24x^2-264x
---
24x^2-264x-11*34848 = 0
x = 132
=================
Cheers,
Stan H.